A record on a turntable is spinning and coming to rest. Att = 0s the angular velocity of the record is 20.0rad / s with a constant acceleration df - 5.00rad / (s ^ 2) What the record's angular velocity at t = 0.500 s?


1
Expert's answer
2021-02-25T11:25:07-0500

At t = 0 s,

Angular velocity, "\omega_o=20rad/s"

Angular acceleration, "\alpha = 5rad/s^2"

At t = 0.500 s,

"\omega=\omega_o+\alpha t"

"\Rightarrow\omega=20+5\times0.500"

"\Rightarrow \omega=22.5rad/s"